OpenStudy (rayman5766):

Solve for x 2x+3/12=2x/8

OpenStudy (campbell_st):

is the question \[\frac{2x + 3}{12} = \frac{2x}{8}\]

OpenStudy (rayman5766):

yes sir

OpenStudy (rayman5766):

I think it is -1/7 but it only lets me typ in one thing so

OpenStudy (campbell_st):

cross multiply to remove the denominators \[8(2x + 3) = 12(2x)\] distribute the 8 on the left and it becomes 16x + 24 = 24x can you solve it from here...?

OpenStudy (rayman5766):

yes hold on stay here

OpenStudy (rayman5766):

its 3

OpenStudy (campbell_st):

that's correct, well done

OpenStudy (rayman5766):

thanks I try